# Low-Overhead Channel Estimation via 3D Extrapolation for TDD mmWave Massive MIMO Systems Under High-Mobility Scenarios

### v1

@@ Abstract
In time division duplexing (TDD) millimeter wave (mmWave) massive multiple-input multiple-output (MIMO) systems, the downlink channel state information (CSI) can be attained through uplink channel estimation thanks to the uplink-downlink channel reciprocity. However, the channel aging issue is significant under high-mobility scenarios and thus necessitates frequent uplink channel estimation. In addition, large amounts of antennas and subcarriers lead to high-dimensional CSI matrices, aggravating the pilot training overhead. To systematically reduce the pilot overhead, a spatial, frequency, and temporal domain (3D) channel extrapolation framework is proposed in this paper. Considering the marginal effects of pilots in the spatial and frequency domains and the effectiveness of traditional knowledge-driven channel estimation methods, we first propose a knowledge-and-data driven spatial-frequency channel extrapolation network (KDD-SFCEN) for uplink channel estimation by exploiting the least square estimator for coarse channel estimation and joint spatial-frequency channel extrapolation to reduce the spatial-frequency domain pilot overhead. Particularly, we propose the attention-based sub-element extrapolation module and the progressive extrapolation architecture to improve the accuracy of joint spatial-frequency channel extrapolation. Then, resorting to the uplink-downlink channel reciprocity and temporal domain dependencies of downlink channels, a temporal uplink-downlink channel extrapolation network (TUDCEN) is proposed for slot-level channel extrapolation, aiming to enlarge the pilot signal period and thus reduce the temporal domain pilot overhead under high-mobility scenarios. Specifically, we propose the spatial-frequency sampling embedding module to reduce the representation dimension and consequent computational complexity, and we propose to exploit the autoregressive generative Transformer for generating downlink channels autoregressively thanks to the powerful capability of generative artificial intelligence. Numerical results demonstrate the superiority of the proposed framework in significantly reducing the pilot training overhead by more than $16$ times and improving the system’s spectral efficiency under high-mobility scenarios.
@@
Channel Extrapolation, High-Mobility Scenarios, Millimeter Wave, Uplink-Downlink Channel Reciprocity, Massive MIMO
@@ I Introduction
MILLIMETER wave (mmWave) massive multi-input multi-output (MIMO) has been recognized as a pivotal technology for the fifth generation (5G) wireless communication systems and beyond [1, 2, 3]. mmWave massive MIMO is anticipated to provide spatial degrees of freedom, diversity or multiplexing gain, and array gain, thereby improving the spectral and energy efficiencies of wireless communication systems. To reap the benefits of massive MIMO, accurate channel state information (CSI) should be attained whether operating in the time division duplexing (TDD) or frequency division duplexing (FDD) modes[4, 5]. In FDD massive MIMO systems, the base station (BS) needs to obtain the downlink CSI through downlink channel estimation and CSI feedback from the user equipment (UE) due to the lack of uplink-downlink channel reciprocity. While in TDD massive MIMO systems, thanks to the uplink-downlink channel reciprocity, the BS can obtain the downlink CSI via uplink channel estimation and derive the downlink CSI from the estimated uplink CSI.
The uplink-downlink channel reciprocity in the TDD massive MIMO systems holds within the channel coherence time. However, in high-mobility scenarios, the channel is fast time-varying due to the UE movement, resulting in a short channel coherence time and the consequent channel aging issue, i.e., the channel varies between when it is acquired at the BS and when it is used for downlink precoding[6]. In addition, compared with sub-6 GHz bands, mmWave bands are more vulnerable to the UE movement since higher frequency bands generally lead to a shorter channel coherence time[7], deteriorating the channel aging issue in mmWave massive MIMO systems under high-mobility scenarios. To avoid significant performance degradation caused by channel aging, frequent channel estimation needs to be conducted such that the downlink precoding can be conducted based on up-to-date CSI, leading to huge temporal domain pilot training overhead. Besides, due to the large number of antennas at the BS and large amounts of subcarriers in orthogonal frequency-division multiplexing (OFDM) systems, the spatial and frequency domain pilot training overhead is also huge and unacceptable. In addition, to reduce the hardware cost and power consumption, hybrid beamforming structure is usually adopted and only a small number of radio frequency (RF) chains are deployed at the BS, especially for a millimeter wave (mmWave) massive MIMO system. To obtain the uplink CSI at all receiving antennas at the BS, the BS has to switch the antennas connected to these RF chains several times during the uplink channel estimation, leading to substantial time and power consumption. Therefore, it is of great significance to minimize the pilot training overhead in the spatial, frequency, and temporal domains while estimating the CSI accurately.
@@
In addition to spatial and frequency domain channel extrapolation, channel prediction (i.e., temporal domain channel extrapolation) is also a convincing method for alleviating the huge pilot training overhead in high-mobility scenarios. Linear extrapolation methods [14] and statistical prediction models, e.g., autoregressive (AR) models [15, 16], were proposed and have demonstrated that channel prediction is promising to mitigate the impacts of channel aging. The advancements of DNNs further improve the channel prediction performance and thus would further reduce the channel estimation frequency in high-mobility scenarios. In [17], a novel long short-term memory (LSTM) based channel predictor was proposed to learn channel variations and thereby reduce the pilot training overhead for channel estimation. An attention-based channel predictor was proposed in [18] to achieve frame-level channel estimates for mobile scenarios. Nonetheless, existing channel prediction methods are unable to deal with massive MIMO-OFDM systems with substantial antennas and subcarriers due to the high-dimensional CSI matrices and the consequent tremendous computational complexity. In addition, since the channel aging issue is critical in high-mobility scenarios, frame-level channel prediction methods are unable to achieve high spectral efficiency due to the varying channels, posing demands on slot-level channel extrapolation. However, existing channel prediction methods predict future channels at the same time granularity as the collected historical channels. This indicates that exploiting these methods for slot-level channel extrapolation comes at the cost of sacrificing a large amount of time-frequency resources for slot-level historical channel estimation, making them infeasible in practice.
@@
To overcome the limitations of these existing works and to systematically reduce the pilot overhead, a spatial, frequency, and temporal domain (3D) channel extrapolation framework is proposed in this paper to reduce the pilot training overhead from these three domains respectively. First, it can be observed that the number of pilots in one certain domain shows the marginal effect, i.e., as the number of pilots in this domain increases, the improvement in channel estimation accuracy gradually decreases. [^1] This observation is further validated with our simulation results in Section V-B. Due to the capability of deep neural networks (DNNs) in extracting the spatial and frequency domain characteristics of massive MIMO channels, it is expected that exploiting DNNs for joint spatial and frequency domain channel extrapolation will further reduce the pilot training overhead. However, although some works have pointed out that jointly learning spatial-frequency domain features is beneficial to channel estimation[11, 19], spatial-frequency channel extrapolation has not yet been well investigated to reduce the pilot training overhead for mmWave massive MIMO-OFDM systems. Therefore, considering the marginal effects of pilots in the spatial and frequency domains and the effectiveness of traditional knowledge-driven channel estimation methods, we first propose a knowledge-and-data driven spatial-frequency channel extrapolation network (KDD-SFCEN) for uplink channel estimation by exploiting the least square estimator for coarse channel estimation and joint spatial-frequency channel extrapolation to reduce the spatial-frequency domain pilot overhead. Then, resorting to the uplink-downlink channel reciprocity and temporal domain dependencies of downlink channels, a temporal uplink-downlink channel extrapolation network (TUDCEN) is proposed for slot-level channel extrapolation, aiming to enlarge the pilot signal period and thus reduce the temporal domain pilot overhead. To the best of our knowledge, none of the existing works have investigated such a systematic framework to reduce the pilot training overhead and improve the spectral efficiency of mmWave massive MIMO-OFDM systems under high-mobility scenarios. The major contributions of this paper are summarized as follows:
@@
    We propose the KDD-SFCEN to reduce the spatial-frequency domain pilot overhead effectively via joint spatial-frequency channel extrapolation. In addition, it is worth emphasizing that the substantial time and power consumption due to antenna switching can be avoided through spatial extrapolation. The proposed KDD-SFCEN consists of a knowledge-driven coarse channel estimator to provide coarse channel estimates and accelerate the training of the proposed network. The KDD-SFCEN also comprises a spatial-frequency channel extrapolator encompassing the attention-based sub-element extrapolation module and the progressive extrapolation architecture. The proposed attention-based sub-element extrapolation module mimics the sub-pixel imaging technology and is demonstrated to be highly effective in learning spatial-frequency channel characteristics for joint spatial and frequency extrapolation. Moreover, the progressive extrapolation architecture is proposed to progressively extrapolate the uplink CSI, thereby further improving the performance of joint spatial-frequency channel extrapolation with few pilots in the spatial and frequency domains.
@@
    We propose the TUDCEN for accurate slot-level channel extrapolation given the estimated uplink CSI at the first slot. Through the proposed TUDCEN, channel estimation can be conducted less frequently and more slots can be configured for data transmission under high-mobility scenarios, further reducing the pilot training overhead from the temporal domain perspective and improving the system’s spectral efficiency. The TUDCEN is composed of an uplink-downlink channel calibration network (UDCCN) and a downlink channel extrapolation network (DCEN). The UDCCN calibrates the estimated uplink channel to the downlink channel at the first downlink slot to compensate for the hardware asymmetry in transceivers. The proposed DCEN achieves slot-level channel extrapolation via spatial-frequency sampling embedding and autoregressive generation with the generative Transformer[20]. Spatial-frequency sampling embedding is proposed to reduce the tremendous computational complexity due to high-dimensional CSI matrices. Specifically, the spatial-frequency sampling embedding layer samples the spatial and frequency domain representation of downlink channels with a spatial sampling factor and a frequency sampling factor to reduce the representation dimension by resorting to the spatial and frequency correlations of downlink channels. Meanwhile, the sampled antenna groups and subcarrier groups are combined to facilitate model training. We exploit the generative Transformer for generating downlink channels autoregressively thanks to the powerful capability of generative artificial intelligence (AI), which is able to conduct slot-level downlink channel extrapolation given only the first downlink channel, thereby avoiding heavy slot-level historical downlink channel estimation.
@@
    We use the sounding reference signal (SRS) defined by the 3rd generation partnership project (3GPP) 5G technical specification [21] as the pilot signal for uplink pilot training. The frame structures and system settings in framework design and numerical simulations also follow the 3GPP 5G technical specification. Numerical results demonstrate the superiority of the proposed framework in significantly reducing the spatial-frequency domain pilot overhead by more than $4$ times via spatial-frequency channel extrapolation. In addition, via enlarging the pilot signal period with slot-level channel extrapolation, the proposed framework further reduces the temporal domain pilot overhead by $4$ times and significantly improves the mmWave massive MIMO system’s spectral efficiency under high-mobility scenarios.
The remainder of this paper is organized as follows. In Section II, we introduce the system model and formulate the spatial, frequency, and temporal channel extrapolation problem. In Section III, we propose the KDD-SFCEN for uplink channel estimation to reduce the spatial-frequency domain pilot training overhead. In Section IV, we propose the TUDCEN for slot-level channel extrapolation to enlarge the pilot signal period and thus reduce the temporal domain pilot overhead. In Section V, simulation results are presented to demonstrate the superiority of the proposed spatial, frequency, and temporal channel extrapolation framework. Finally, we conclude this work in Section VI.
@@ II-A System Model
As shown in Fig. 1 (a), we consider a mmWave massive MIMO system working in the TDD mode, where a single BS equipped with $N_{T}\gg 1$ antennas and $N_{RF}\ll N_{T}$ RF chains serves a single user equipped with $N_{R}$ antennas. The BS is realized with a hybrid precoding architecture, as Fig. 1 (b) shows. The system operates with the OFDM modulation with a total of $N_{c}\gg 1$ subcarriers. By leveraging the uplink-downlink channel reciprocity in the TDD system, the downlink channel can be acquired via uplink channel estimation to avoid downlink channel estimation and feedback at the UE side, thereby improving the system’s spectral efficiency. A simplified downlink CSI acquisition and data transmission process based on uplink pilot training is also shown in Fig. 1 (a). The process consists of three phases, i.e., uplink pilot transmission (phase $1$), BS signal processing (phase $2$), and downlink data transmission (phase $3$). Specifically, the uplink pilot signal is first transmitted by the UE to the BS for uplink channel estimation (phase $1$). Then the BS estimates the uplink channel given the received pilot signal, derives the downlink channel, and conducts downlink precoding (phase $2$). After that, the BS transmits data symbols to the UE based on the downlink precoder and other transmission parameters (phase $3$).
@@
Given $\mathbf{H}^{d}_{i}$, the downlink precoding matrix $\mathbf{F}_{i}\in\mathbb{C}^{N_{T}\times N_{s}}$ at the $i$-th subcarrier with $N_{s}$ being the number of data streams and $N_{s}\leq N_{R}$ can be designed by the singular value decomposition (SVD)-based precoding algorithm [23]. Specifically, SVD for $\mathbf{H}^{d}_{i}$ can be expressed as
$\mathbf{H}^{d}_{i}=\mathbf{U}_{i}\boldsymbol{\Sigma}_{i}\mathbf{V}^{H}_{i},$ (6)
where $\mathbf{U}_{i}\in\mathbb{C}^{N_{R}\times N_{R}}$, $\boldsymbol{\Sigma}_{i}\in\mathbb{C}^{N_{R}\times N_{T}}$, and $\mathbf{V}^{H}_{i}\in\mathbb{C}^{N_{T}\times N_{T}}$ are the left singular vectors matrix, the diagonal matrix of singular values, and the right singular vectors matrix, respectively. Note that $\mathbf{F}_{i}=[\mathbf{f}_{i}^{(1)},\mathbf{f}_{i}^{(2)},\ldots,\mathbf{f}_{i}^{(N_{s})}]$, where $\mathbf{f}_{i}^{(j)}\in\mathbb{C}^{N_{T}\times 1}$, $j=1,2,\ldots,N_{s}$ are the first $N_{s}$ column vectors of $\mathbf{V}$. Then, in the downlink transmission, the received signal $\mathbf{y}_{i}\in\mathbb{C}^{N_{R}\times 1}$ at the $i$-th subcarrier can be represented as
$\mathbf{y}_{i}=\mathbf{H}^{d}_{i}\mathbf{F}_{i}\mathbf{s}_{i}+\mathbf{n}_{i},$ (7)
where $\mathbf{s}_{i}\in\mathbb{C}^{N_{s}\times 1}$ and $\mathbf{n}_{i}\in\mathbb{C}^{N_{s}\times 1}$ are the transmitted signal and the downlink noise at the $i$-th subcarrier, respectively.
Figure 2: The overall framework of the proposed spatial, frequency, and temporal channel extrapolation method. The proposed method first realizes uplink channel estimation with the KDD-SFCEN to reduce the spatial-frequency domain pilot training overhead $C_{sl}$, and then conducts accurate slot-level channel extrapolation to enlarge the pilot signal period $T_{p}$ with the TUDCEN, thereby systematically reducing the pilot training overhead $C_{o}$.

### v2

@@ Abstract
In time division duplexing (TDD) millimeter wave (mmWave) massive multiple-input multiple-output (MIMO) systems, downlink channel state information (CSI) can be obtained from uplink channel estimation thanks to channel reciprocity. However, under high-mobility scenarios, frequent uplink channel estimation is needed due to channel aging. Additionally, large amounts of antennas and subcarriers result in high-dimensional CSI matrices, aggravating pilot training overhead. To address this, we propose a three-domain (3D) channel extrapolation framework across spatial, frequency, and temporal domains. First, considering the effectiveness of traditional knowledge-driven channel estimation methods and the marginal effects of pilots in the spatial and frequency domains, a knowledge-and-data driven spatial-frequency channel extrapolation network (KDD-SFCEN) is proposed for uplink channel estimation via joint spatial-frequency channel extrapolation to reduce spatial-frequency domain pilot overhead. Then, leveraging channel reciprocity and temporal dependencies, we propose a temporal uplink-downlink channel extrapolation network (TUDCEN) powered by generative artificial intelligence for slot-level channel extrapolation, aiming to reduce the tremendous temporal domain pilot overhead caused by high mobility. Numerical results demonstrate the superiority of the proposed framework in significantly reducing the pilot training overhead by $16$ times and improving the system’s spectral efficiency under high-mobility scenarios compared with state-of-the-art channel estimation/extrapolation methods.
@@
Channel Extrapolation, Three-Domain, High-Mobility Scenarios, Millimeter Wave, Massive MIMO
@@ I Introduction
MILLIMETER wave (mmWave) massive multi-input multi-output (MIMO) has been recognized as a pivotal technology for the fifth generation (5G) wireless communication systems and beyond [1, 2, 3]. mmWave massive MIMO is anticipated to provide spatial degrees of freedom, diversity or multiplexing gain, and array gain, thereby improving the spectral and energy efficiencies of wireless communication systems. To reap the benefits of massive MIMO, accurate channel state information (CSI) should be attained whether operating in the time division duplexing (TDD) or frequency division duplexing (FDD) modes[4, 5]. Specifically, in TDD massive MIMO systems, thanks to the uplink-downlink channel reciprocity, the base station (BS) can obtain the downlink CSI via uplink channel estimation and downlink CSI derivation at the BS side.
The uplink-downlink channel reciprocity in the TDD massive MIMO systems holds within the channel coherence time. However, in high-mobility scenarios, the channel is fast time-varying due to the UE movement, resulting in a short channel coherence time and the consequent channel aging issue, i.e., the channel varies between when it is acquired at the BS and when it is used for downlink precoding[6]. In addition, compared with sub-6 GHz bands, mmWave bands are more vulnerable to the mobility of the user equipment (UE) since higher frequency bands generally lead to a shorter channel coherence time[7], deteriorating the channel aging issue in mmWave massive MIMO systems under high-mobility scenarios. To avoid significant performance degradation caused by channel aging, frequent channel estimation needs to be conducted such that downlink precoding can be conducted based on up-to-date CSI, leading to huge temporal domain pilot training overhead. Besides, due to the large number of antennas at the BS and large amounts of subcarriers in orthogonal frequency-division multiplexing (OFDM) systems, the spatial and frequency domain pilot training overhead is also huge and unacceptable. In addition, to reduce the hardware cost and power consumption, a hybrid precoding structure is usually adopted and only a small number of radio frequency (RF) chains are deployed at the BS, especially for mmWave massive MIMO systems. To obtain the uplink CSI at all receiving antennas at the BS, the BS has to switch the antennas connected to these RF chains several times during the uplink channel estimation, leading to substantial time and power consumption. Therefore, it is of great significance to minimize the pilot training overhead in the spatial, frequency, and temporal domains while estimating the CSI accurately.
@@
In addition to spatial and frequency domain channel extrapolation, channel prediction (i.e., temporal channel extrapolation) is also a convincing method for alleviating the huge pilot training overhead in high-mobility scenarios. Linear extrapolation methods [14] and statistical prediction models, e.g., autoregressive (AR) models [15, 16], were proposed and have demonstrated that channel prediction is promising to mitigate the impacts of channel aging. The advancements of DNNs further improve the channel prediction performance and thus would further reduce the channel estimation frequency in high-mobility scenarios. In [17], a novel long short-term memory (LSTM) based channel predictor was proposed to learn channel variations and thereby reduce the pilot training overhead for channel estimation. An attention-based channel predictor was proposed in [18] to achieve frame-level channel estimates for mobile scenarios. However, frame-level channel prediction methods are unable to achieve high spectral efficiency due to fast time-varying channels under high-mobility scenarios, posing demands on slot-level channel extrapolation. Moreover, existing channel prediction methods are unable to deal with massive MIMO-OFDM systems with substantial antennas and subcarriers due to the high-dimensional CSI matrices and the consequent tremendous computational complexity.
@@
To overcome the limitations of these existing works and to systematically reduce the pilot overhead, a spatial, frequency, and temporal domain (3D) channel extrapolation framework is proposed in this paper to reduce the pilot training overhead from these three domains respectively. First, it can be observed that the number of pilots in one certain domain shows the marginal effect, i.e., as the number of pilots in this domain increases, the improvement in channel estimation accuracy gradually decreases. [^1] This observation is further validated with our simulation results in Section V-B. Second, considering the capability of deep neural networks (DNNs) in extracting the spatial and frequency domain characteristics of massive MIMO channels, it is expected that exploiting DNNs for joint spatial and frequency domain channel extrapolation will further reduce the pilot training overhead. However, although some works have pointed out that jointly learning spatial-frequency domain features is beneficial to channel estimation[11, 19], spatial-frequency channel extrapolation has not yet been well investigated to reduce the pilot training overhead for mmWave massive MIMO-OFDM systems. In addition, spatial-frequency correlations and long-term temporal domain dependencies (i.e., dependencies over multiple slots) of downlink channels are rarely investigated to reduce computational complexity and facilitate slot-level channel extrapolation. Therefore, we first propose a knowledge-and-data driven spatial-frequency channel extrapolation network (KDD-SFCEN) to reduce the spatial-frequency domain pilot overhead. Specifically, coarse low-dimensional channel estimates are first obtained with a knowledge-driven channel estimator from low-density pilots in spatial and frequency domains and then extrapolated to the full-dimensional CSI with the spatial-frequency channel extrapolation network. After that, a temporal uplink-downlink channel extrapolation network (TUDCEN) is proposed for slot-level channel extrapolation, aiming to reduce the temporal domain pilot overhead. To the best of our knowledge, none of the existing works have investigated such a systematic framework to reduce the pilot training overhead and improve the spectral efficiency of mmWave massive MIMO-OFDM systems under high-mobility scenarios. The major contributions of this paper are summarized as follows:
@@
    We propose the KDD-SFCEN to reduce the spatial-frequency domain pilot overhead effectively via joint spatial-frequency channel extrapolation. In addition, it is worth emphasizing that the substantial time and power consumption due to antenna switching can be avoided through spatial extrapolation. The proposed KDD-SFCEN consists of a knowledge-driven coarse channel estimator to provide coarse channel estimates and accelerate the training of the proposed network. The KDD-SFCEN also comprises a spatial-frequency channel extrapolator encompassing the attention-based sub-element extrapolation module (ASEEM) and the progressive extrapolation architecture to exploit spatial and frequency correlations of wireless channels for spatial and frequency domain extrapolation. Specifically, based on the observation that antennas/subcarriers exhibit stronger correlations when they are closer, the progressive extrapolation architecture is designed to first focus on capturing the correlations between neighboring antennas/subcarriers, and then progressively broaden the scope to capture correlations, while the ASEEM is designed to capture specific spatial/frequency correlations in the process.
@@
    We propose the TUDCEN for accurate slot-level channel extrapolation given the estimated uplink CSI at the first slot. Through the proposed TUDCEN-aided pilot training scheme, channel estimation can be conducted less frequently and more slots can be configured for data transmission than a general pilot training scheme under high-mobility scenarios, further reducing the temporal domain pilot overhead and improving the system’s spectral efficiency. The TUDCEN is composed of an uplink-downlink channel calibration network (UDCCN) and a downlink channel extrapolation network (DCEN). The UDCCN calibrates the estimated uplink channel to the downlink channel at the first downlink slot to compensate for the hardware asymmetry in transceivers[20]. The DCEN achieves slot-level channel extrapolation via spatial-frequency sampling embedding and autoregressive generation with the generative Transformer[21]. Spatial-frequency sampling embedding is proposed to reduce the tremendous computational complexity due to high-dimensional CSI matrices by exploiting spatial and frequency correlations of wireless channels. We exploit the generative Transformer for generating downlink channels autoregressive by learning temporal dependencies of wireless channels over time, which enables slot-level downlink channel extrapolation given only the first downlink channel, thereby avoiding heavy slot-level historical downlink channel estimation.
@@
    We use the sounding reference signal (SRS) defined by the 3rd generation partnership project (3GPP) 5G technical specification [22] as the pilot signal for uplink pilot training. The frame structures and system settings in framework design and numerical simulations also follow the 3GPP 5G technical specification. Numerical results demonstrate the superiority of the proposed framework in significantly reducing the spatial-frequency domain pilot overhead by $4$ times via spatial-frequency channel extrapolation compared with state-of-the-art spatial-frequency channel estimation/extrapolation methods. In addition, by reducing the times of single-slot channel estimations with slot-level channel extrapolation, the proposed framework further reduces the temporal domain pilot overhead by $4$ times and significantly improves the mmWave massive MIMO system’s spectral efficiency under high-mobility scenarios.
The remainder of this paper is organized as follows. In Section II, we introduce the system model and formulate the spatial, frequency, and temporal channel extrapolation problem. In Section III, we propose the KDD-SFCEN for uplink channel estimation to reduce the spatial-frequency domain pilot training overhead. In Section IV, we propose the TUDCEN for slot-level channel extrapolation to reduce the temporal domain pilot overhead. In Section V, simulation results are presented to demonstrate the superiority of the proposed spatial, frequency, and temporal channel extrapolation framework. Finally, we conclude this work in Section VI.
@@ II-A System Model
As shown in Fig. 1 (a), we consider a mmWave massive MIMO system working in the TDD mode, where a single BS equipped with $N_{T}\gg 1$ antennas and $N_{RF}\ll N_{T}$ RF chains serves a single user equipped with $N_{R}$ antennas. The BS is realized with a hybrid precoding architecture, as Fig. 1 (b) shows. The system operates with the OFDM modulation with a total of $N_{c}\gg 1$ subcarriers. By leveraging the uplink-downlink channel reciprocity in the TDD system, the downlink CSI can be acquired via uplink channel estimation and downlink channel derivation. A simplified downlink CSI acquisition and data transmission process based on uplink pilot training is also shown in Fig. 1 (a). The process consists of three phases, i.e., uplink pilot transmission (phase $1$), BS signal processing (phase $2$), and downlink data transmission (phase $3$). Specifically, the uplink pilot signal is first transmitted by the UE to the BS for uplink channel estimation (phase $1$). Then the BS estimates the uplink channel given the received pilot signal, derives the downlink channel, and conducts downlink precoding (phase $2$). After that, the BS transmits data symbols to the UE based on the downlink precoder and other transmission parameters (phase $3$). [^2] Note that the working modes of the hybrid array in phase $1$ and phase $3$ are different. In phase $1$, each RF chain is connected to an antenna at a time to receive the individual signal from this antenna and switches to connect to other antennas, which enables CSI acquisition for all antennas and facilitates downlink precoding. While in phase $3$, each RF chain is connected to multiple antennas at the same time and combines the signals received by multiple antennas, achieving a balance between system performance and complexity.
@@
After that, hybrid precoding is conducted for downlink transmission and thus the received signal $\mathbf{y}_{i}\in\mathbb{C}^{N_{R}\times 1}$ at the $i$-th subcarrier can be represented as
$\mathbf{y}_{i}=\mathbf{H}^{d}_{i}\mathbf{F}^{RF}_{i}\mathbf{F}^{BB}_{i}\mathbf{s}_{i}+\mathbf{n}_{i},$ (6)
where $\mathbf{s}_{i}\in\mathbb{C}^{N_{R}\times 1}$ and $\mathbf{n}_{i}\in\mathbb{C}^{N_{R}\times 1}$ are the transmitted signal and the downlink noise at the $i$-th subcarrier, and $\mathbf{F}^{RF}_{i}\in\mathbb{C}^{N_{T}\times N_{RF}}$ and $\mathbf{F}^{BB}_{i}\in\mathbb{C}^{N_{RF}\times N_{R}}$ are the analog and digital precoding matrices, respectively. In addition, $\mathbf{F}^{RF}_{i}$ can be designed via the discrete Fourier transform (DFT) codebook-based analog precoding algorithm, and $\mathbf{F}^{BB}_{i}$ can be designed via the zero-forcing (ZF) algorithm.
Figure 2: The overall framework of the proposed spatial, frequency, and temporal channel extrapolation method. The proposed method first realizes uplink channel estimation with the KDD-SFCEN to reduce the spatial-frequency domain pilot training overhead $C_{sl}$, and then conducts accurate slot-level channel extrapolation with the TUDCEN to reduce the times of uplink channel estimations $\frac{T}{T_{p}}$, thereby systematically reducing the pilot training overhead $C_{o}$.